This is from a game show, would like to know how to set it up, been YEARS since I did word problems. I know you assign X to what's missing, but this one is more complicated
Tom was asked his age. He replied, " In 2 years, I will be twice as old as I was 5 years ago." How old is he?
X would be his age now, X + 2 would be his age in two years and I am lost after that except there is a 2x and a -5 in there somewhere. PLease help! THank you!

Answer by Theo(13342)

x = your age now
x + 2 = your age in 2 years
x - 5 = your age 5 years ago.

In 2 years, I will be twice as old as I was 5 years ago." How old is he?

your equation is x + 2 = 5 * (x-5)

solve that equation for x.

you will get x = 12

x + 2 = 14
x - 5 = 7
14 is 2 * 7